dria:

Pn=Pn-1 + 30, with initial population P0=200. a. Calculate P1 and P2 b. Find an explicit formula for Pn c. Use your formula to predict the number of houses in 10 years d. When will the number of houses reach 400 houses?

jhonyy9:

so just substitute inside given formula what you know a. p_1 = p1-1 +30 = p0 +30 = ? p_2 = p2-1 +30 = p_1 +30 = ?